OpenStudy (anonymous):

A 2.80kg mass is dropped from a height of 4.50m. Find the PE and KE when it is 3.00m above the ground.

OpenStudy (anonymous):

Initial PE = 2.8 * 10 * 4.5

OpenStudy (anonymous):

PE at 3m height = 2.8 * 10 * 4.5 KE at 3m height = Change in PE = 2.8*10(4.5-3)

OpenStudy (anonymous):

@_loree22

OpenStudy (anonymous):

My PE is correct but my KE isn't. I got 42 but when I plug the number in, its mark as wrong.

OpenStudy (anonymous):

@vikrantg4

OpenStudy (anonymous):

PE at 3 m height = 2.8 * 10 * 3 = 84 , Sorry My bad KE is 42 !

OpenStudy (anonymous):

Thank you :) @vikrantg4

OpenStudy (anonymous):

Anytime !

OpenStudy (anonymous):

Can you help me find the height for this problem. A 0.155kg arrow is shot upward at 31.4m/s. what is its PE.

OpenStudy (anonymous):

calculate the kinetic energy first.. the same energy will become PE at max height

OpenStudy (anonymous):

I did, I got 76.4119. When I plugged the number in I got 116.0696761.

OpenStudy (anonymous):

Put it keeps saying I got it wrong, what did I do wrong ?

OpenStudy (anonymous):

What is KE from one of these values ?

OpenStudy (anonymous):

My KE is 76.411 and for PE I got 116.0696761.

OpenStudy (anonymous):

No No.. The KE is 76.411 This must be the PE at max height according to conservation of mechanical energy. So PE = mgh = 76.411 h = 76.411/ (9.8 * 0.155) = 50.3 m

OpenStudy (anonymous):

Thank you :)